OpenStudy (anonymous):

What is the square root of -72?

OpenStudy (anonymous):

You can't take the square root of a negative number.

OpenStudy (anonymous):

Unless you are dealing with imaginary number

OpenStudy (anonymous):

If your asking for the solutions to $$x^2=-72$$ They are $$-i\sqrt{72},i\sqrt{72}$$

OpenStudy (psymon):

Assuming you need imaginary solutions then yes, the square root of -72 would just be: \[i \sqrt{72}\]. But of course that can be simplified further ifyou know how to :3

OpenStudy (anonymous):

well, some of my answer choices are imaginary. -6 sqrt2 6sqrt-2 6sqrt2i 6isqrt2

OpenStudy (anonymous):

\[\sqrt{-72}=i \sqrt{72}= i \sqrt{9*8}= i \sqrt{9}\sqrt{8}=3i \sqrt{8}=3i \sqrt{2*4}=6i \sqrt{2}\] Do you understand the process of doing that?